1. Given the following equation:

_____ K2PtCl4 + _____ NH3 --------> _____ Pt(NH3)2Cl2 + _____ KCl

a) Balance the chemical equation.

b) Determine the theoretical yield of KCl if you start with 34.5 grams of NH3.
1
Expert's answer
2017-05-25T11:29:10-0400
K2PtCl4 + 2 NH3 --------> Pt(NH3)2Cl2 + 2 KCl

M(NH3)=14+3=17 g/mol
n(NH3)=m/M=34.5/17=2.03 mol
n(NH3)=n(KCl) <according to the balance>
M(KCl)=39+35.5=74.5 g/mol
m(KCl)=n*M=2.03*74.5=151.2 g